# SkewRoute: Training-Free LLM Routing for Knowledge Graph Retrieval- Augmented Generation via Score Skewness of Retrieved Context

### v1

@@ Abstract
Large language models excel at many tasks but often incur high inference costs during deployment. To mitigate hallucination, many systems use a knowledge graph to enhance retrieval-augmented generation (KG-RAG). However, the large amount of retrieved knowledge contexts increase these inference costs further. A promising solution to balance performance and cost is LLM routing, which directs simple queries to smaller LLMs and complex ones to larger LLMs. However, no dedicated routing methods currently exist for RAG, and existing training-based routers face challenges scaling to this domain due to the need for extensive training data. We observe that the score distributions produced by the retrieval scorer strongly correlate with query difficulty. Based on this, we propose a novel, training-free routing framework—the first tailored to KG-RAG—that effectively balances performance and cost in a plug-and-play manner. Experiments show our method reduces calls to larger LLMs by up to 50% without sacrificing response quality, demonstrating its potential for efficient and scalable LLM deployment.
@@
Hairu Wang1,3thanks: Equal contribution., Yuan Feng1,311footnotemark: 1, Yukun Cao1,3, Xike Xie2,3thanks: Corresponding author., S Kevin Zhou2,3 1School of Computer Science, University of Science and Technology of China, China 2School of Biomedical Engineering, USTC, China 3Data Darkness Lab, MIRACLE Center, Suzhou Institute for Advanced Reasearch, USTC, China {hrwang00,yfung}@mail.ustc.edu.cn, {xkxie,skevinzhou}@ustc.edu.cn
@@ 1 Introduction
To reduce LLM inference costs in applications, recent studies have focused on LLM routing Aggarwal et al. (2024); Ding et al. (2024); Ong et al. (2025); Feng et al. (2025), routing simple questions to smaller LLMs and complex questions to larger LLMs. This is primarily because the performance benefits of increased model scale exhibit diminishing returns; for example, Qwen2.5-14B-Instruct incurs less than twice the inference cost of Qwen2.5-7b-Instruct while delivering a 7.45% performance increase. In contrast, Qwen2.5-72b-Instruct requires almost 7x the inference cost of Qwen2.5-14b-Instruct but only achieves a 2.12% improvement, as illustrated in Figure 2(b). However, while recent routing methods have been explored in direct LLM QA applications, their application and extensibility to RAG domains is limited, primarily due to two reasons:
@@
    To the best of our knowledge, this work is the first to introduce a training-free routing framework tailored for KG-RAG, leveraging the skewness patterns in retrieved scores to dynamically assess query difficulty, offering a novel perspective for future efficient RAG research.
@@ 2 Preliminaries
Power-law. Power-law Song et al. (2024); Havrilla and Liao (2024), ubiquitous in nature, reveals an inverse relationship between scores and their descending ranks: $S(n)\;=\;\frac{C}{n^{\alpha}}$. In this formulation, $n$ denotes the rank and $S(n)$ is the corresponding score. In log-log space, $logS(n)$ against $log(n)$ approximates a straight line with slope $-\alpha$. While in linear space, it appears as a few high scores followed by a long tail of lower ones. In KG-RAG, we observe that for certain queries, the scores of retrieved contexts also follow this power-law behavior.
KG-RAG. RAG integrates contexts retrieved from external knowledge bases to enhance the accuracy and credibility of LLMs. In KG-RAG, the external knowledge base is a KG, and the retrieved contexts take the form of triples $\{\tau=(h,r,t)\}$. Existing research on KG-RAG generally involves two stages: retrieval and generation. For example, the state-of-the-art KG-RAG method, SubgraphRAG trains a scorer $\mathcal{R}$ to select a subset of triples $T=\{\tau_{1},\tau_{2},\dots,\tau_{K}\}$ by their scores $W=\{s_{1},s_{2},...,s_{K}\}$ relevant to $q$ from $G$ in the retrieval phase. And then feeds $q$, an instruction $I$, and $T$ to a generator $\mathcal{G}$ (typically an LLM) to produce the final answer: $\mathcal{G}(I,T,q)\rightarrow a$, where $a\in\mathcal{A}$. Knowledge Graph Question Answering (KGQA) is often used to evaluate KG-RAG, aiming to obtain answers from KGs given query $q$, where $q\in Q$.
@@ 3.1 Skewness of Retrieved Context Scores
Figure 3 presents two typical examples of the top-100 context score distributions for different queries, showing clear and significant differences. Power-law distribution is highly skewed with long-tail. In Figure 3(a) and Figure 3(c), the distribution exhibits power-law behavior and shows power-law behavior: only a few contexts are relevant to the question, while most score below 0.1 and are irrelevant. By contrast, Figure 3(b) and Figure 3(d) show lower skewness, with a large portion of contexts clearly relevant—100% contexts have scores above 0.1. This difference in skewness intuitively corresponds well with the query difficulties: difficult queries naturally require reasoning across more knowledge contexts, resulting in lower skewness.
3.2 Correlation Between Score Skewness and Query Difficulty
Retrieved context scores vary in skewness across queries. To measure it, a natural metric is the area under the scores. Highly skewed distributions, where scores drop off rapidly and the low-score long tail dominates, results in smaller areas. In contrast, low skewness distributions with a wealth of high-score triples and a gentle initial descent, yield a much larger area under the scores. Concretely, we first min-max normalize the raw scores to $[0,1]$, eliminating the magnitude differences of raw scores across different queries. And then sum these normalized scores to approximate the area. Using this metric, the area computed in Figure 3(c) is 1.07, whereas in Figure 3(d) it is 65.65.
In KG-RAG, the retrieval quality directly impacts the answer generation of LLMs. To assess query difficulty, we use a straightforward metric: we check whether the top-ranked context in retrieved set contains the answer for given query. Simple query requires only a single knowledge context to answer. If the Top-$1$ contains real answer, we consider this query is simple. In this case, the scorer has accurately identified the context critical for reasoning with high confidence.
Next, we investigate the relationship between score skewness of retrieved contexts and query difficulty. To this end, we extract queries whose answers appear in the top-scoring context and compute the areas under their scores. Smaller areas correspond to higher skewness. We partition the area from their minimum to maximum into several contiguous intervals, each corresponding to a distinct level of skewness. Figure 12 illustrates the count of simple queries across different skewness levels. Overall, a consistent trend was observed in both datasets, which queries are predominantly concentrated at higher skewness levels. This allows us to select simple queries based on their skewness. The finding reveals the strong correlation between query difficulty and score skewness, which can serve as effective indicator for routing LLMs to balance performance and cost. Further validation results is provided in the Appendix A.3.3.
@@ 3.2 Correlation Between Score Skewness and Query Difficulty
Figure 4: Simple Query Number Across Skewness.
3.3 Using Skewness to Route
@@ 9      Add $ans$ to $\mathcal{A}$;
In this section, we present our training-free routing framework, which can be seamlessly integrated into KG-RAG in a plug-and-play manner, as shown in Algorithm 1. By leveraging the strong correlation between score skewness of retrieved contexts and query difficulty, we require no additional training. And any metric for measuring distribution skewness is compatible with our framework for LLM selection. Queries with high score skewness are routed to economical smaller LLMs, while those with low score skewness are directed to more powerful larger LLMs, thereby striking an optimal balance between response quality and inference cost. As previously mentioned, area under scores can serve as a metric of distribution skewness, yet it still has important drawbacks. First, it is highly sensitive to min-max normalization: differences in raw score ranges across queries lead to inconsistent scaling and unstable area comparisons. Second, it collapses the entire scores into a single value, sacrificing crucial information about its distribution shape. The consequence is that widely divergent distributions can produce similar area, undermining its ability to differentiate distinct skewness. To address these shortcomings, we further explore new solutions to measure score skewness. We propose a cumulative threshold-based routing method and analyze scores from the perspective of entropy and gini coefficient to make routing decisions. Compared with the straightforward area under scores, these methods offer superior and robust discriminative power between simple and difficult queries.
@@
Cumulative Threshold-based Routing In probability and statistics, the Cumulative Distribution Function (CDF) Kolmogoroff (1933); Kolmogorov (2018) quantifies the probability that a random variable $X$ is less than or equal to a value $x$, offering a comprehensive view of the distribution. Queries with high score skewness due to the rapid decay rate, require only a few contexts to reach a cumulative probability. Conversely, queries with low score skewness necessitate a larger number of contexts to achieve the same value. Thus, cumulative threshold-based routing could guide LLM selection. The contexts need to rank in a descending order and then normalize by the score sum into a probability distribution: $p_{i}=s_{i}/\sum_{j=1}^{K}s_{j}$. Building upon this, the cumulative sum $C_{k}$ of them is computed by: $C_{k}=\sum_{i=1}^{k}p_{i}$. For a given probability $P$ (e.g., 95%), we find the smallest $k$ satisfying $C_{k}\geq P$. Queries are classified as simple if $k\leq\theta_{k}$, otherwise difficult.
@@
Entropy-based Routing Entropy Shannon (1948) can quantify the uniformity of a probability distribution. Queries with low score skewness, characterized by with more uniform and slower decaying context scores, exhibit higher entropy, while highly skewed distributions correspond to lower entropy. Thus, entropy can effectively assess query difficulty based on context scores, enabling dynamic routing of queries. Calculating the entropy for scores requires converting them into a probability distribution $p=\{p_{1},p_{2},\dots,p_{K}\}$ by normalizing the scores $p_{i}=s_{i}/\sum_{j=1}^{K}s_{j}$. The entropy is then calculated as $H=-\sum_{i=1}^{N}p_{i}\log_{2}(p_{i})$. Thereby, a query is routed as simple if the entropy of context scores falls below a predetermined threshold $\theta_{H}$, and as difficult otherwise.
@@
Gini Coefficient-based Routing The Gini coefficient Gates (1909); Lorenz (1905), traditionally measuring income inequality, can effectively reflect the imbalance of context scores. A higher Gini coefficient indicates greater skewness of context scores, reflecting a more pronounced imbalance among the retrieved knowledge contexts. Thus, Gini coefficient can serve as an effective metric for routing decision. The calculation of the Gini coefficient for context scores $W$ involves sorting them in ascending order $s^{\prime}_{1}\leq s^{\prime}_{2}\leq\dots\leq s^{\prime}_{K}$ and applying the following formula: $\frac{1}{K}\left(K+1-2\frac{\sum_{i=1}^{K}(K-i+1)s^{\prime}_{i}}{\sum_{j=1}^{K}s^{\prime}_{j}}\right)$. Queries are routed as simple if the calculated Gini coefficient exceeds a threshold $\theta_{G}$, indicating high skewness, and as difficult if it falls below $\theta_{G}$.
@@ 4.1 Settings
(a) Qwen Hit@1
(b) Qwen F1-Score
(c) Llama Hit@1
(d) Llama F1-Score
Figure 5: Results of Routing on WebQSP. (a)(b) between Qwen7b/72b and (c)(d) between Llama8b/70b
(a) Qwen Hit@1
(b) Qwen F1-Score
(c) Llama Hit@1
(d) Llama F1-Score
Figure 6: Results of Routing on CWQ. (a)(b) between Qwen7b/72b and (c)(d) between Llama8b/70b
Datasets. To test the effectiveness of routing on knowledge-intensive reasoning tasks, we used two widely recognized KGQA datasets: CWQ Talmor and Berant (2018) and WebQSP Yih et al. (2016). Freebase Bollacker et al. (2008) serves as the background KG for both datasets, comprising more than 120 million triples. Query difficulty, assessed by the number of required knowledge contexts, varies across datasets. Specifically, WebQSP is relatively simple, containing queries with only 1 or 2 knowledge contexts. In contrast, CWQ contains more difficult reasoning problems involving up to 4 contexts. Further details of the datasets are described in Appendix A.1.
Baselines. SubgraphRAG, the latest SOTA method in KG-RAG, achieves effective and efficient retrieval of a suitable amount of relevant contexts from KGs for LLM generation. Consequently, all experiments conducted in this study are based on it. We consider three baselines: all-larger-LLM strategy, all-smaller-LLM, and random mixing.
Reasoning LLMs. We utilize a diverse selection of open-source models for our experiments: Qwen2.5-7b-Instruct, Qwen2.5-72b-Instruct Qwen et al. (2025), Llama3.1-8b-Instruct, and Llama3.1-70b-Instruct Llama Team (2024). For brevity, we will refer to them as Qwen7b, Qwen72b, Llama8b, and Llama70b, respectively. Based on their parameter counts, Qwen7b and Llama8b are classified as smaller LLMs, while their counterparts, Qwen72b and Llama70b, fall under the category of larger LLMs. Generally, larger LLMs with more parameters, deliver stronger inference performance but incur greater overhead.
Evaluation Metrics. Consistent with prior research, we employ Hit@1  [^2] Here, we use the corrected evaluation code of SubgraphRAG. In the paper, the original Hit@1 of Llama3.1-8b-Instruct with 100 triples is 56.98. and F1-Score as evaluation metrics. Hit@1 measures the rate at which the correct answer appears at the top prediction position. Given that a query may have multiple answers, F1-Score is used to assess the comprehensive coverage of all correct answers, effectively balancing precision and recall. The objective of LLM routing is to balance performance and cost, so the final evaluation metric is the performance at a specified larger LLM call ratio, equivalently Hit@1 or F1-Score under a given inference cost.
Experiment Implementations. For reasoning, we use one-shot prompting to LLMs’ generation. Consistent with SubgraphRAG, the prompt template we employe is detailed in Appendix A.4. The temperature parameter is set to 0 for reproduction, and the maximum token length for generation is fixed at 4000.

### v2

@@ Abstract
Large language models excel at many tasks but often incur high inference costs during deployment. To mitigate hallucination, many systems use a knowledge graph to enhance retrieval-augmented generation (KG-RAG). However, the large amount of retrieved knowledge contexts increase these inference costs further. A promising solution to balance performance and cost is LLM routing, which directs simple queries to smaller LLMs and complex ones to larger LLMs. However, no dedicated routing methods currently exist for RAG, and existing training-based routers face challenges scaling to this domain due to the need for extensive training data. We observe that the score distributions produced by the retrieval scorer strongly correlate with query difficulty. Based on this, we propose an extremely simple yet effective routing framework, the first specifically designed for KG-RAG that efficiently balances performance and cost in a plug-and-play manner. It delivers over 3x higher routing effectiveness while reducing runtime to less than 0.001x compared to existing methods. Our code is available at https://github.com/hrwang00/SkewRoute.
@@
Hairu Wang1,3thanks: Equal contribution., Yuan Feng1,311footnotemark: 1, Yukun Cao1,3, Xike Xie2,3thanks: Corresponding author., S Kevin Zhou2,3 1School of Computer Science, University of Science and Technology of China, China 2School of Biomedical Engineering, USTC, China 3Suzhou Institute for Advanced Reasearch, USTC, China {hrwang00,yfung}@mail.ustc.edu.cn, {xkxie,skevinzhou}@ustc.edu.cn
@@ 1 Introduction
To reduce LLM inference costs in applications, recent studies have focused on LLM routing Aggarwal et al. (2024); Ding et al. (2024); Ong et al. (2025); Feng et al. (2025), routing simple questions to smaller LLMs and complex questions to larger LLMs. This is primarily because the performance benefits of increased model scale exhibit diminishing returns; for example, Qwen2.5-14B-Instruct incurs less than twice the inference cost of Qwen2.5-7b-Instruct while delivering a 7.45% performance increase. In contrast, Qwen2.5-72b-Instruct requires almost 7x the inference cost of Qwen2.5-14b but only achieves a 2.12% improvement, as illustrated in Figure 2(b). However, while recent routing methods have been explored in direct LLM QA applications, their application and extensibility to RAG domains is limited, primarily due to two reasons:
@@
    To the best of our knowledge, this work is the first to introduce a training-free routing framework tailored for KG-RAG, leveraging the skewness patterns in retrieved scores to dynamically assess query difficulty, offering a novel perspective for future efficient RAG research
@@ 2 Preliminaries
Power-law. Power-law Song et al. (2024); Havrilla and Liao (2024), ubiquitous in nature, reveals an inverse relationship between scores and their descending ranks: $S(n)\;=\;\frac{C}{n^{\alpha}}$. In this formulation, $n$ denotes the rank and $S(n)$ is the corresponding score. In log-log space, $logS(n)$ against $log(n)$ approximates a straight line with slope $\alpha$. While in linear space, it appears as a few high scores followed by a long tail of lower ones. In KG-RAG, we observe that for certain queries, the scores of retrieved contexts also follow this power-law behavior.
KG-RAG. RAG integrates contexts retrieved from external knowledge bases to enhance LLMs. In KG-RAG, the knowledge bases are KGs, and the retrieved contexts take the form of triples $\{\tau=(h,r,t)\}$. Existing research on KG-RAG generally involves two stages: retrieval and generation. For example, the state-of-the-art KG-RAG method, SubgraphRAG trains a scorer $\mathcal{R}$ to select a subset of triples $T=\{\tau_{1},\tau_{2},\dots,\tau_{K}\}$ by their scores $W=\{s_{1},s_{2},\dots,s_{K}\}$ relevant to $q$ from $G$ in the retrieval phase. And then feeds $q$, an instruction $I$, and $T$ to a generator $\mathcal{G}$ (typically an LLM) to produce the final answer: $\mathcal{G}(I,T,q)\rightarrow a$, where $a\in\mathcal{A}$. Knowledge Graph Question Answering (KGQA) is often used to evaluate KG-RAG, aiming to obtain answers from KGs given query $q$, where $q\in Q$.
@@ 3.1 Skewness of Retrieved Context Scores
Figure 3 presents two typical examples of the top-100 context score distributions for different queries, showing clear and significant differences. Power-law distribution is highly skewed with long-tail. In Figure 3(a) and Figure 3(c), the distribution exhibits power-law behavior and approximates a straight line at the loglog scale: only a few contexts are relevant to the question, while most score below 0.1 and are irrelevant. By contrast, Figure 3(b) and Figure 3(d) show lower skewness, with a large portion of contexts clearly relevant—100% contexts have scores above 0.1. This difference in skewness intuitively corresponds well with the query difficulty: difficult queries naturally require reasoning across more knowledge contexts, resulting in lower skewness.
3.2 Correlation Between Skewness and Query Difficulty
We directly quantify skewness using the area under the min-max normalized score curve: a smaller area indicates high skewness (rapid score drop-off), while a larger area reflects low skewness (many high scores and gradual decline). For example, Figure 3(c) has an area of 1.07, whereas Figure 3(d) has 65.65.
However, query difficulty does not have explicit labels for direct assessment. Moreover, in RAG, retrieval context directly affects difficulty: clear and sufficient contexts make questions inherently easier. Therefore, we approximate question difficulty by analyzing the scorer’s retrieval results. In an extreme case, if the top-ranked triple context already contains the answer, the question is relatively easy, as it can be answered by matching a single fact from the KG. The scorer identifies the correct triple with high confidence during retrieval. Conversely, if the answer-containing triple context appears lower in the ranked list, this typically indicates a more difficult, multi-hop reasoning scenario, where several semantically related but indirect triples are retrieved. Thus, we use the rank of the answer-containing triple as a proxy for query difficulty: a lower rank (i.e., further down the list) corresponds to higher difficulty.
Using the above metrics, we examine the correlation between the skewness of retrieved scores and query difficulty. Queries are divided into three equally sized groups based on the area size, where a smaller area indicates higher skewness. As shown in Figure 4, higher skewness is strongly associated with lower query difficulty, as the answer-containing context is more likely to be top-ranked. This insight allows us to classify queries as simple or difficult based on score skewness and, accordingly, route them to LLMs of different scales to balance inference cost and performance.  [^2] Further analysis is provided in Appendix A.7.
@@ 3.2 Correlation Between Skewness and Query Difficulty
Figure 4: Query Difficulty Across Score Skewness.
@@ 3.3 Using Skewness to Route
In this section, we present our training-free routing framework, which can be seamlessly integrated into KG-RAG in a plug-and-play manner, as shown in Algorithm 1. By leveraging the strong correlation between score skewness of retrieved contexts and query difficulty, we require no additional training. And any metric for measuring distribution skewness is compatible with our framework for LLM selection. Queries with high score skewness are routed to economical smaller LLMs, while those with low skewness are directed to more powerful larger LLMs, thereby striking an optimal balance between response quality and inference cost.
As previously mentioned, area under scores can serve as a metric of distribution skewness, yet it still has important drawbacks. First, it is highly sensitive to min-max normalization: differences in raw score ranges across queries lead to inconsistent scaling and unstable area comparisons. Second, it collapses the entire scores into a single value, sacrificing crucial information about its distribution shape. The consequence is that widely divergent distributions can produce similar area, undermining its ability to differentiate distinct skewness. To address these shortcomings, we further explore new solutions to measure score skewness. We propose a cumulative threshold-based routing method and analyze scores from the perspective of entropy and gini coefficient to make routing decisions. Compared with the straightforward area under scores, these methods offer superior and robust discriminative power between simple and difficult queries.
@@
Cumulative Threshold-based Routing. In probability and statistics, Cumulative Distribution Function Kolmogoroff (1933); Kolmogorov (2018) quantifies the probability a random variable $X$ is less than or equal to value $x$, offering a comprehensive view of the distribution. Queries with high score skewness due to the rapid decay, require only a few contexts to reach a cumulative probability. Conversely, queries with low skewness need more contexts to achieve the same value. Thus, cumulative threshold-based routing could guide LLM selection. The context scores $W$ need to rank in a descending order and then normalize by the score sum into a probability distribution: $p_{i}=s_{i}/\sum_{j=1}^{K}s_{j}$. Building upon this, the cumulative sum $C_{k}$ of them is computed by: $C_{k}=\sum_{i=1}^{k}p_{i}$. For a given probability $P$ (e.g., 95%), we find the smallest $k$ satisfying $C_{k}\geq P$. The function $\mathcal{L}(W,\theta)$ is defined as $k-\theta_{k}$. Queries are classified as simple if $k\leq\theta_{k}$, otherwise difficult.
@@
Entropy-based Routing Entropy Shannon (1948) can quantify the uniformity of a probability distribution. Queries with low score skewness, characterized by with more uniform and slower decaying context scores, exhibit higher entropy, while highly skewed distributions correspond to lower entropy. Thus, entropy can effectively assess query difficulty based on context scores, enabling dynamic routing of queries. Calculating the entropy for scores $W$ requires converting them into a probability distribution $p=\{p_{1},p_{2},\dots,p_{K}\}$ by normalizing the scores $p_{i}=s_{i}/\sum_{j=1}^{K}s_{j}$. The entropy is then calculated as $H=-\sum_{i=1}^{N}p_{i}\log_{2}(p_{i})$. In this context, $\mathcal{L}(W,\theta)$ refers to $H-\theta_{H}$. A query is routed as simple if the entropy of context scores falls below a predetermined threshold $\theta_{H}$, and as difficult otherwise.
@@
Gini Coefficient-based Routing The Gini coefficient Gates (1909); Lorenz (1905), traditionally measuring income inequality, can effectively reflect the imbalance of context scores. A higher Gini coefficient indicates greater skewness of context scores, reflecting a more pronounced imbalance among the retrieved knowledge contexts. Thus, Gini coefficient can serve as an effective metric for routing decision. The calculation of the Gini coefficient for context scores $W$ involves sorting them in ascending order $s^{\prime}_{1}\leq s^{\prime}_{2}\leq\dots\leq s^{\prime}_{K}$ and applying the following formula: $Gini=\frac{1}{K}\left(K+1-2\frac{\sum_{i=1}^{K}(K-i+1)s^{\prime}_{i}}{\sum_{j=1}^{K}s^{\prime}_{j}}\right)$. Here, $\mathcal{L}(W,\theta)$ is $\theta_{G}-Gini$. Queries are routed as simple if the calculated Gini coefficient exceeds a threshold $\theta_{G}$, indicating high skewness, and as difficult if it falls below $\theta_{G}$.
@@ 4.1 Settings
Datasets. To test the effectiveness of routing on knowledge-intensive reasoning tasks, we used two widely recognized KGQA datasets: CWQ Talmor and Berant (2018) and WebQSP Yih et al. (2016). Freebase Bollacker et al. (2008) serves as the KG for both datasets, comprising more than 120 million triples. Query difficulty, assessed by the number of required contexts, varies across datasets. Specifically, WebQSP is relatively simple, containing queries with only 1 or 2 knowledge contexts. Conversely, CWQ contains more difficult reasoning problems involving up to 4 contexts. Further details are described in Appendix A.1.
Baselines. SubgraphRAG, the latest SOTA method in KG-RAG, achieves effective and efficient performance. We thereby adopt its scorers in this study. We compare with the following baselines: Random Routing, which serves as the lower bound. Under the same budget, improvements over this baseline demonstrate the effectiveness of other routing methods. RouteLLM Ong et al. (2025) and GraphRouter Feng et al. (2025) are included as representatives of the latest routing methods. Following their claim of generalization across unseen data and LLMs, we use their pre-trained weights as the routing baseline against our training-free framework. Although these methods involve additional training cost, they still face significant challenges under the RAG setting.
Reasoning LLMs. We utilize a diverse selection of open-source models for our experiments: Qwen2.5-7B-Instruct, Qwen2.5-72B-Instruct Qwen et al. (2025), Llama3.1-8B-Instruct, and Llama3.1-70B-Instruct Llama Team (2024). Based on their parameter counts, Qwen7B and Llama8B are classified as smaller LLMs, while their counterparts, Qwen72B and Llama70B, fall under the larger LLMs.
Evaluation Metrics. Consistent with prior research, we employ Hit@1  [^3] Here, we use the officially updated evaluation code of SubgraphRAG, which leads to slightly lower results compared to those reported in the original paper (see https://github.com/Graph-COM/SubgraphRAG). However, this does not affect the validity of our routing method’s evaluation. as evaluation metrics. It measures the rate at which the correct answer appears at the top prediction position. The objective of LLM routing is to balance performance and cost, which means a higher Hit@1 is preferable under the same call ratio of the larger LLM. Thereby, we define Average Effectiveness (Avg. Eff.) as the mean improvement in Hit@1 over random routing.